OpenStudy (anonymous):

Evaluate 8 to the 2 thirds power.

OpenStudy (anonymous):

8^(2/3)=4. use a calculator

mathslover (mathslover):

Instead of using a calculator , you can solve it like this : \(8^{\cfrac{2}{3}} = \sf{\sqrt[3]{8^2}} \\ \sf{\sqrt[3]{8\times 8}\\ \sqrt[3]{(4\times 2)\times (4\times 2)}\\ \sqrt[3]{4\times 4\times 4} = 4 }\)

mathslover (mathslover):

Also, welcome to OpenStudy @albo97 !

OpenStudy (anonymous):

Thank you for the help

mathslover (mathslover):

You're welcome :) Also, if you think, I help you then you can award me with a medal by clicking on the "Best Response" button .